OpenStudy (anonymous):

condense the expression to a single logarithm: 3logx-4logy-log(x-4)

OpenStudy (nincompoop):

minus means division

OpenStudy (anonymous):

i've gotten it down to logx^3/logy^4 divided by log(x-4) but i don't know how to condense it farther than that, any thoughts?

OpenStudy (nincompoop):

x-4 would be?

OpenStudy (anonymous):

what do you mean?

OpenStudy (anonymous):

sorry still don't understand. I'm not sure what to do with the last -log(x-4)

OpenStudy (anonymous):

\[\log(a)+\log(b)-\log(c) = \log(\frac{ab}{c})\] Should be the last step you need :)

OpenStudy (anonymous):

even though the logarithm has three minus and no addition?

OpenStudy (anonymous):

I showed what to do with addition and subtraction. If it's all subtraction, then just do that.